Ticket Number: HEA-NOC/20020112-01 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> UCC
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020112 03:40
Ticket Closed: Problem End:
OSPF flaps between leabhair.hea.net and UCC access router.
All traffic to and from UCC.
20021012 11:00 BN
Contacted Denis O'Sullivan and informed him of problem. Connection back for 7 hours and seems internal to UCC.
20020113 17:07 BN
Connection flapping again, OSPF flapping between 16:08 and 16:45. However while connection seems to have come back after that ping times very, very slow. Contacted Denis O'Sullivan in UCC, he is going on site to investigate. Further investigations by HEAnet NOC point to excessive CPU utilisation.
20020113 20:00 BN
Denis O'Sullivan rebooted Cork Access router OSPF back up finally at 18:28. Ping times still very slow. Ticket remains open.
20020114 23:15 OSPF flaps on Cork PoP access circuit for 2 minutes.
20020114 05:05 OSPF peering down for 16 seconds.
20020114 19:53 BN
OSPF flaps up and down until 20:34.
20020115 12:24 BN
OSPF flaps again between 12:08 and 12:09. Contacted Denis O'Sullivan in UCC, they are working on problem.
14:23 BN 12:08 - 12:09 OSPF flaps caused by unrelated router reboot.
20020116 00:38-00:54 OSPF flaps cause loss of connectivity.
20020116 03:27 OSPF flap for less than 1 minute
20020116 09:45 BN
Denis O'Sullivan reports further problems in UCC resulting in major access problems to the UCC Lan. These are being investigated by UCC. However intermitent outages are likely.
20020117 07:55 UCC access router power cycled at 17:54, further brief
outages at 21:36, 21:51, 23:43, 23:53, 23:56, 23:58, 00:02
20020117 12:30 BN
Problem seems to be due to attempted attack on UCC network. This is being actively pursued by UCC computing services.
20020117 15:41 BN
UCC reports flaps and connectivity issues seemingly due to SYN attacks on httpd. Connection has been stable today, but ticket
will be kept open. UCC have sent an email regarding the attacks
to JAnet Cert.
20020119 5:35 AB Cicuit continues to flap thoughout the day of 19th. Approx 60 up/down states within 24 hours.
Time to Fix:
Unknown Additional Notes:
Ticket being kept open for monitoring.